OpenStudy (flvskidd):

Let f(x) = 3x2 – x + 2 and g(x) = 5x2 – 1. Find f(g(x)). Show each step of your work.

OpenStudy (anonymous):

ok can u plz show what u have tried already?

OpenStudy (dayakar):

f(x) = 3x^2 -x+2--------(1) g(x) = 5x2-1----------(2) f(g(x) = f[ 5x2-1] [ from(2) g(x) = 5x2-1]

OpenStudy (dayakar):

is it clear

OpenStudy (flvskidd):

sorry i didnt show a pic

OpenStudy (flvskidd):

yes I undertsand

OpenStudy (dayakar):

in equation (1) replace x with 5x2 -1 and simplify it

OpenStudy (dayakar):

f(5x2-1) =3[5x2-1]^2 -[5x2-1] +2 simplify it

OpenStudy (flvskidd):

3(5x-1^2-1)^2-x+2

OpenStudy (flvskidd):

nvm okay

OpenStudy (dayakar):

can u simplify

OpenStudy (flvskidd):

yeah

OpenStudy (flvskidd):

75x^4-35x^2+6

OpenStudy (flvskidd):

?

jhonyy9 (jhonyy9):

@FLVSKidd do you can writing your calculi step by step here please ?

jhonyy9 (jhonyy9):

like (5x^2 -1)^2 = ?

jhonyy9 (jhonyy9):

so there you have wrote ,,show each step of your work" - so than make it please

jhonyy9 (jhonyy9):

ok. (5x^2 -1)^2 = 25x^4 -10x^2 +1 - so this is the first step ,can you continuing please ?

OpenStudy (flvskidd):

did I calculate right? dayakar

OpenStudy (flvskidd):

@dayakar

OpenStudy (dayakar):

plz , wait

jhonyy9 (jhonyy9):

75x^4-35x^2+6 yes - this is right sure, -- but you not have wrote here step by step

OpenStudy (dayakar):

yes, corrct

jhonyy9 (jhonyy9):

hope so much that you understand it now what is the way to calculi like exercise yes ,

OpenStudy (dayakar):

@FLVSKidd .CORRECT